# Hierarchical Mamba Meets Hyperbolic Geometry: A New Paradigm for Structured Language Embeddings

### v2

@@ A.2 Mamba2
This paper introduces a Mamba2-based LLM known as SentenceMamba-16M, a lightweight model with 16M parameters, suitable for resource-efficient and high-quality sentence embedding generation. As we can see in Fig. 1, we incorporate 4 Mamba2 blocks in our SentenceMamba-16M for efficient state-space modeling.
Appendix B Hyperbolic Loss Calculations
The losses can be either triplet loss or contrastive loss as shown in Fig. B1. Depending on our dataset, we can either apply Triplet Loss when we have a triplet relationship in our data and need to enforce relative distance constraints or apply Contrastive Loss when we have pairwise relationships in our data and need to classify pairs as similar or dissimilar.
Fig. B1 illustrates how Hyperbolic Losses are calculated. Based on either triplet relationship data or pairwise relationship data, we perform triplet loss or contrastive loss calculation. Then, we calculate the weighted loss of centripetal loss and contrastive loss as the hyperbolic loss. Loss minimization involves clustering related entities and distancing unrelated ones (clustering loss) and tightening parent entities closer to the hyperbolic manifold’s origin than their child counterparts (centripetal loss), giving it a hierarchical structure.
Figure B1: Calculation of hyperbolic loss from clustering loss and centripetal loss.
Appendix C Dataset Statistics
To provide a comprehensive overview of the datasets used in our experiments, we detail the size, number of entities (nodes), and train/validation/test splits for each dataset in Table C4. These datasets are represented as directed acyclic graphs (DAGs), where nodes denote entities (e.g., diseases in DOID, synsets in WordNet) and edges denote direct subsumption relations (is-a). Splits are created by sampling direct ($E$) and indirect (multi-hop, $T$) subsumptions, ensuring coverage of both mixed-hop prediction and multi-hop inference tasks.
Table C4: Statistics of hierarchical ontology datasets
@@ Appendix C Dataset Statistics
Appendix D Task Formulations
D.1 Multi-Hop inference
@@ D.2 Mixed-Hop prediction
Appendix E Performance Comparisons between Hyperbolic embeddings and Euclidean embeddings
@@
For mixed-hop prediction (Figure E1), HiM-Lorentz achieves better performance on datasets with deeper hierarchies, such as DOID ($\delta$-hyperbolicity = 0.019) and SNOMED-CT ($\delta$-hyperbolicity = 0.026). This aligns with the Lorentz manifold’s ability to handle deeply nested structures more effectively. However, for FoodOn ($\delta$-hyperbolicity = 0.185), HiM-Poincaré slightly outperforms HiM-Lorentz. FoodOn’s higher $\delta$-hyperbolicity indicates a less tree-like structure, suggesting that the Poincaré model’s bounded nature may better capture less hierarchical relationships in certain contexts. Both hyperbolic models significantly outperform the Euclidean baselines.
@@
In multi-hop inference (Figure E2), HiM-Lorentz again demonstrates robust performance, particularly on SNOMED-CT and WordNet, which exhibit deeper hierarchies (SNOMED-CT $\delta$-hyperbolicity = 0.0254, WordNet $\delta$-hyperbolicity = 0.1431). The smaller standard deviations in HiM-Lorentz’s metrics (e.g., 0.003 for SNOMED-CT F1) compared to HiM-Poincaré (0.028) highlight its stability, a benefit of the Lorentz manifold’s numerical advantages. Notably, HiM-Poincaré achieves a slightly higher recall on SNOMED-CT, suggesting that the bounded nature of the Poincaré ball can occasionally enhance sensitivity. However, the overall F1 score favors HiM-Lorentz, indicating better balance in precision and recall.
@@
Figures E3 to E5 illustrate the training dynamics of HiM-Poincaré and HiM-Lorentz across epochs for each dataset and task, plotting Hyperbolic Loss and F1 Score. The hyperbolic loss decreases steadily for both models across all datasets, indicating effective optimization of hierarchical relationships. HiM-Lorentz often exhibits a slightly faster convergence rate and lower final loss compared to HiM-Poincaré, reflecting the Lorentz manifold’s suitability for capturing exponential hierarchical expansion. The F1 Score trends mirror the loss behavior, with HiM-Lorentz often achieving slightly better F1 scores.
Figure E1: Comparisons of mixed-hop prediction performance for DOID, FoodOn, WordNet, and SNOMED-CT datasets based on our proposed hyperbolic mamba (HiM) models–HiM-Poincaré and HiM-Lorentz, and their Euclidean counterparts–Pretrained and fine-tuned sentenceMamba-16M.
Figure E2: Comparisons of multi-hop inference performance for DOID, FoodOn, WordNet, and SNOMED-CT datasets based on our proposed hyperbolic mamba (HiM) models–HiM-Poincaré and HiM-Lorentz, and their Euclidean counterparts–Pretrained and fine-tuned sentenceMamba-16M.
@@
Figure E3: Comparison of hyperbolic loss and F1 score on DOID mixed-hop prediction across epochs.
@@
Figure E4: Comparison of hyperbolic loss and F1 score on WordNet multi-hop inference across epochs.
@@
Figure E5: Comparison of hyperbolic loss and F1 score on SNOMED-CT mixed-hop prediction across the epochs.
Appendix F Code Availability

### v3

@@ A.1.2 Mamba2
This paper introduces a Mamba2-based LLM known as SentenceMamba-16M, a lightweight model with 16M parameters, suitable for resource-efficient and high-quality sentence embedding generation. As we can see in Figure 1, we incorporate 4 Mamba2 blocks in our SentenceMamba-16M for efficient state-space modeling.
A.2 Hyperbolic Loss Calculations
Depending on our dataset, we can either apply Triplet Loss when we have a triplet relationship in our data and need to enforce relative distance constraints or apply Contrastive Loss when we have pairwise relationships in our data and need to classify pairs as similar or dissimilar. Based on either triplet relationship data or pairwise relationship data, we perform triplet loss or contrastive loss calculation. Then, we calculate the weighted loss of centripetal loss and clustering loss as the hyperbolic loss. Hyperbolic loss computation framework is illustrated in Figure 4, which shows how these two loss components are weighted and combined to create the final training objective. Loss minimization involves clustering related entities and distancing unrelated ones (clustering loss) and tightening parent entities closer to the hyperbolic manifold’s origin than their child counterparts (centripetal loss), giving it a hierarchical structure.
Figure 4: Calculation of hyperbolic loss from clustering loss and centripetal loss.
A.3 Dataset Statistics
To provide a comprehensive overview of the datasets used in our experiments, we detail the size, number of entities (nodes), and train/validation/test splits for each dataset in Table 3. These datasets are represented as directed acyclic graphs (DAGs), where nodes denote entities (e.g., diseases in DOID, synsets in WordNet) and edges denote direct subsumption relations (is-a). Splits are created by sampling direct ($E$) and indirect (multi-hop, $T$) subsumptions, ensuring coverage of both mixed-hop prediction and multi-hop inference tasks.
Table 3: Statistics of hierarchical ontology datasets
@@ A.3 Dataset Statistics
A.4 Task Formulations
A.4.1 Multi-Hop inference
@@ A.4.2 Mixed-Hop prediction
A.5 Learning Interpretable Hierarchical Semantics Through Hyperbolic Geometry
To provide more interpretable results of our HiM models for the hierarchical learning, we conducted a deeper geometric analysis of the hyperbolic entity embeddings for semantically related WordNet entities under HiM-Poincaré and HiM-Lorentz manifolds (see the visualization of hyperbolic embeddings in Figure  2). Specifically, we computed three key metrics with the learned hyperbolic embeddings: 1) “hyperbolic geodesic distances” between each pair of entities, 2) “h-norm” represents the norm distance from the origin, a higher h-norm often indicates a deeper or more specific concept in the hierarchy, 3) “depth” is the WordNet tree depth. In both sets of entities, the h-norm correlates strongly with the hierarchical depth, see Figure 5. For instance, in Table 4, the entity sport (depth 9) has an h-norm of 0.55, while skateboarding (depth 12) has an h-norm of 2.25, reflecting the hierarchical expansion from general to specific concepts. However, a key difference emerges when comparing the two manifolds: HiM-Lorentz consistently produces smaller hyperbolic distances and h-norms compared to HiM-Poincaré. For example, in Table 5 (HiM-Lorentz), the parent-child relationships (e.g., sport $\rightarrow$ skating $\rightarrow$ skateboarding) have tighter distances and h-norm gradients compared to Table 4.
This reduction in distances under the Lorentz manifold is advantageous for hierarchical modeling. The Lorentz model’s unbounded nature avoids the boundary constraints of the Poincaré ball, which can lead to numerical instability near the boundary. By mapping embeddings into an unbounded hyperboloid, HiM-Lorentz achieves tighter clustering of related entities (e.g., between sport and skating: 0.73 h-Norm of HiM-Lorentz vs. 1.67 h-norm of HiM-Poincaré) while maintaining the hierarchical structure. This tighter clustering enhances the model’s ability to distinguish fine-grained relationships, especially in deeper hierarchies, as evidenced by the smaller standard deviations of HiM-Lorentz in performance metrics (Table 1).
Table 4: Hyperbolic distances, h-norms, and depths for sports-related entities (from Figure 2) using HiM-Poincaré, sorted by increasing depth.
sport outdoor sport skating rowing skateboarding speed skating
sport 0.00 1.17 1.67 1.62 2.43 2.40
outdoor sport 1.17 0.00 1.90 1.95 2.66 2.62
skating 1.67 1.90 0.00 2.36 3.12 3.07
rowing 1.62 1.95 2.36 0.00 3.10 3.11
skateboarding 2.43 2.66 3.12 3.10 0.00 3.76
speed skating 2.40 2.62 3.07 3.11 3.76 0.00
h-norm 0.55 0.97 1.48 1.48 2.25 2.25
depth 9 10 11 11 12 12
Table 5: Hyperbolic distances, h-norms, and depths for sports-related entities (from Figure 2) using HiM-Lorentz, sorted by increasing depth.
sport outdoor sport skating rowing skateboarding speed skating
sport 0.00 0.55 0.73 0.71 0.87 0.87
outdoor sport 0.55 0.00 0.79 0.87 0.98 0.95
skating 0.73 0.79 0.00 0.95 1.08 1.08
rowing 0.71 0.87 0.95 0.00 1.07 1.07
skateboarding 0.87 0.98 1.08 1.07 0.00 1.20
speed skating 0.87 0.95 1.08 1.07 1.20 0.00
h-norm 0.30 0.48 0.65 0.65 0.81 0.81
depth 9 10 11 11 12 12
A.6 Performance Comparisons between Hyperbolic embeddings and Euclidean embeddings
@@
For mixed-hop prediction (Figure 6), HiM-Lorentz achieves better performance on datasets with deeper hierarchies, such as DOID ($\delta$-hyperbolicity = 0.019) and SNOMED-CT ($\delta$-hyperbolicity = 0.026). This aligns with the Lorentz manifold’s ability to handle deeply nested structures more effectively. However, for FoodOn ($\delta$-hyperbolicity = 0.185), HiM-Poincaré slightly outperforms HiM-Lorentz. FoodOn’s higher $\delta$-hyperbolicity indicates a less tree-like structure, suggesting that the Poincaré model’s bounded nature may better capture less hierarchical relationships in certain contexts. Both hyperbolic models significantly outperform the Euclidean baselines.
@@
In multi-hop inference (Figure 7), HiM-Lorentz again demonstrates robust performance, particularly on SNOMED-CT and WordNet, which exhibit deeper hierarchies (SNOMED-CT $\delta$-hyperbolicity = 0.0254, WordNet $\delta$-hyperbolicity = 0.1431). The smaller standard deviations in HiM-Lorentz’s metrics (e.g., 0.003 for SNOMED-CT F1) compared to HiM-Poincaré (0.028) highlight its stability, a benefit of the Lorentz manifold’s numerical advantages. Notably, HiM-Poincaré achieves a slightly higher recall on SNOMED-CT, suggesting that the bounded nature of the Poincaré ball can occasionally enhance sensitivity. However, the overall F1 score favors HiM-Lorentz, indicating better balance in precision and recall.
@@
Figure 6: Comparisons of mixed-hop prediction performance for DOID, FoodOn, WordNet, and SNOMED-CT datasets based on our proposed hyperbolic mamba (HiM) models–HiM-Poincaré and HiM-Lorentz, and their Euclidean counterparts–Pretrained and fine-tuned sentenceMamba-16M.
Figure 7: Comparisons of multi-hop inference performance for DOID, FoodOn, WordNet, and SNOMED-CT datasets based on our proposed hyperbolic mamba (HiM) models–HiM-Poincaré and HiM-Lorentz, and their Euclidean counterparts–Pretrained and fine-tuned sentenceMamba-16M.
Figures 8 to 10 illustrate the training dynamics of HiM-Poincaré and HiM-Lorentz across epochs for each dataset and task, plotting Hyperbolic Loss and F1 Score. The Hyperbolic Loss decreases steadily for both models across all datasets, indicating effective optimization of hierarchical relationships. HiM-Lorentz often exhibits a slightly faster convergence rate and lower final loss compared to HiM-Poincaré, reflecting the Lorentz manifold’s suitability for capturing exponential hierarchical expansion. The F1 Score trends mirror the loss behavior, with HiM-Lorentz often achieving slightly better F1 scores.
@@
Figure 8: Comparison of hyperbolic loss and F1 score on DOID mixed-hop prediction across epochs.
@@
Figure 9: Comparison of hyperbolic loss and F1 score on WordNet multi-hop inference across epochs.
@@
Figure 10: Comparison of hyperbolic loss and F1 score on SNOMED-CT mixed-hop prediction across the epochs.
A.7 Additional Experimental Analysis
A.7.1 Prompted LLM Experiment (HiM vs. GPT-4o)
To evaluate HiM’s performance relative to contemporary large language models, we conducted a zero-shot evaluation using GPT-4o [^3] https://openai.com/index/hello-gpt-4o/ on the WordNet mixed-hop prediction task. This comparison provides insight into how our specialized hyperbolic architecture performs against general-purpose language models that rely on vast pretraining but lack explicit hierarchical inductive biases. We generated 500 binary classification questions following the structure “Is [entity1] a subtype/subclass of [entity2]?” sampled from the same test set used for HiM evaluation. The experimental setup mirrored HiM’s training regime: for each sampled child node, we generated one positive question and ten negative questions (corresponding to HiM’s 1 positive parent + 10 hard negatives). GPT-4o was provided with a list of 74,401 WordNet entities as context and answered all 500 questions in a single zero-shot prompt without additional training. The results in Table 6 demonstrate that both HiM variants substantially outperform GPT-4o. GPT-4o performs well on general knowledge hierarchies where concepts like ‘dog’ → ‘animal’ are well-represented in large-scale training corpora. HiM (both Poincaré and Lorentz) still outperform GPT-4o by a clear margin, even on such general-knowledge hierarchy. The superior performance of HiM highlights the effectiveness of our hyperbolic modeling approach for hierarchical reasoning, even when compared to pretrained LLM with significantly larger parameter counts and extensive pretraining.
Table 6: F1 scores comparing HiM models with GPT-4o on WordNet mixed-hop prediction task
Dataset HiM GPT-4o
Poincaré Lorentz
WordNet-mixed 0.859 0.850 0.750
A.7.2 Computational Efficiency Analysis
To substantiate our claims regarding Mamba’s linear complexity advantages, we conducted a sequence length scaling study on the WordNet dataset for the mixed-hop prediction task. The results in Table 7 validate Mamba’s theoretical linear complexity characteristics. Doubling the sequence length from 128 to 256 tokens results in an exact 2× increase in both FLOPs (4.46G → 8.9G) and MACs (2.23G → 4.45G), while memory consumption remains constant at 66.91MB due to the fixed number of model parameters and activations. This linear scaling behavior contrasts sharply with transformer-based architectures, where sequence length increases would result in quadratic growth in computational requirements.
Table 7: Sequence length scaling analysis for HiM-16M-Poincaré on WordNet mixed-hop prediction
Sequence Length FLOPs MACs Memory Training Time
128 4.46 G 2.23 G 66.91 MB 0:41:32
256 8.90 G 4.45 G 66.91 MB 0:42:13
A.8 Code Availability
@@
A.9 Declaration of LLM usage
LLMs were only used to assist with writing and formatting, not as part of the core methodology. Large Language Models (LLMs) were used in a limited capacity during the preparation of this manuscript for grammar checking and text refinement. All technical contributions, results and insights are the original work of the authors.